# Low Photon Number Non-Invasive Imaging Through Time-Varying Diffusers

### v1

@@ Abstract
Scattering poses a significant challenge in optical imaging. In microscopy, it leads to progressive degradation of image quality preventing sample examination at increasing depths. As the thickness of the sample increases, the number of photons that can successfully pass through it decreases, leading to a reduced signal and preventing us from obtaining clear images. Imaging becomes even more difficult with dynamic scatterers like biological tissues that change over time. We demonstrate non-invasive imaging of fluorescent objects hidden behind dynamic diffusers. Our reconstruction uses a stack of images recorded for different diffuser realizations and works even when individual camera frames contain small photon numbers.
keywords:
Imaging in dynamic complex media, Single photon detection, Non-invasive imaging, Speckle pattern analysis
@@ 1 Introduction
Scattering plays a crucial role in determining the appearance of observed objects. When scattering occurs multiple times within a medium, such as biological tissues, fog, or clouds, it is referred to as a complex medium. Light propagation through complex media results in random interference patterns known as speckle patterns[1]. For a long time, light scattering in complex media was considered uncontrollable. This fundamental challenge of light scattering affects not only imaging [2], but also poses obstacles in communication, spectroscopy, and other optical and photonic technologies.
Imaging through complex media has long been seen as a formidable challenge. However, a groundbreaking experiment by Vellekoop and Mosk [3] demonstrated that by controlling wavefronts through spatial light modulators (SLMs), light can be focused through scattering media. Subsequent experiments by Popoff et al. made significant advancements by introducing a method to measure the transmission matrix of complex media, allowing for wavefront control using SLMs behind these media [4, 5].
These remarkable advancements led to a better understanding of complex media and paved the way for developing non-invasive imaging techniques through such media. Specifically, experiments [6, 7] utilized the phenomenon known as the memory effect. Within small regions, when an input wavefront reaching a complex medium is tilted within a certain angular range, the output wavefront is equally tilted, resulting in the translation of the speckle pattern behind the complex medium [8, 9]. In imaging, within a memory effect region, light emitted from each point source behind the complex medium creates an almost identical speckle pattern that is shifted according to the position of the source origin.
@@
While several techniques for imaging in complex media were developed and represent significant achievements, their widespread adoption remains limited. One limitation is that these techniques typically work with static diffusers. In live bioimaging, biological tissues change over time, necessitating methods that can handle dynamic scattering. Similarly, LIDAR must operate in fog, which changes over time. Some approaches to overcome this issue involve using high-speed digital micromirror devices (DMDs) [15, 16, 17, 18, 19] as spatial light modulators (SLMs) for fast transmission matrix measurement. However, all these methods require high signal levels, which hinder their practical applications. For rapidly changing diffusers, one must acquire light over a short period, resulting in low light intensities being captured. This creates a growing need for robust imaging techniques that work in low-light scenarios through dynamic complex media.
Here, we demonstrate a non-invasive technique for imaging behind dynamic scattering layers. This technique is capable of operating in the low-photon regime and with single photon cameras, with much less than one photon per pixel on average . We demonstrate our technique in a widefield fluorescence microscope to image astrocyte cells stained with the common dye Alexa Fluor 568 through a dynamically changing diffuser, recording less than one photon per camera pixel per frame. Our method does not require wavefront shaping, transmission matrix characterization, or light focusing through the complex medium.
@@ 2 Principles
Figure 1: Concept of our method. (a) Schematic of our experiment. We record a movie of a luminescent object behind a changing diffuser. Each frame is assumed to be captured fast enough to treat the diffuser as static during the exposure. (b) Each frame $I_{i}$ is a convolution of the object $O$ with the speckle PSF $S_{i}$ of the varying diffuser. (c) The spatial Fourier transform of (b). (d) Averaging the Fourier transform moduli of all frames approximates the modulus of the Fourier transform of the object image without scattering. (e) The phase of the complex Fourier transform, lost in this process, is retrieved using a phase retrieval algorithm [20].
Our method for imaging through a dynamic diffuser is based on techniques such as those presented in [6, 7, 10]. The method involves recording a movie of a luminescent object behind a changing diffuser (see Figure 1a). We assume that the frame duration is short enough for the diffuser to remain static during each frame. We demonstrate that our approach enables robust reconstruction even if individual frames do not have a sufficient detected photon number for object reconstruction.
When the medium changes during the integration time, the resulting speckle pattern blurs, leading to an averaged-out image that loses spatial information about the sample. To overcome this, rather than directly integrating the signal recorded by the camera, we divide the signal into short-exposure, low-photon frames. We compute the Fourier transform modulus for each frame and average the resulting matrices using root mean square. This method enables effective imaging in dynamic media, yielding improved results, particularly in low-light scenarios, compared to the techniques described by [7, 6, 10].
Since each frame $I_{i}$ captured by the camera is a convolution of the object $O$ with the speckle Point Spread Function (PSF) $S_{i}$ of the varying diffuser (as illustrated in 1b), and the object is within the memory effect range [8, 9] (meaning each point of the object has the same PSF $S_{i}$), we have:
$I_{i}=O*S_{i},$
where $i=1,2,3,\ldots,N$ is the frame index, and $N$ is the total number of frames.
This implies that
$\mathcal{F}\{I_{i}\}=\mathcal{F}\{O\}\mathcal{F}\{S_{i}\},$
@@
where $\mathcal{F}\{I_{i}\}$, $\mathcal{F}\{O\}$, and $\mathcal{F}\{S_{i}\}$ denote the spatial Fourier transforms of each frame $I_{i}$, the object $O$, and the speckle Point Spread Function $S_{i}$, respectively (as shown in 1c). Then,
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}.$
For many realizations of the diffuser, the root mean square of the Fourier transform moduli of the frames for the object recorded through the diffuser converges to the modulus of the Fourier transform of the object image without scattering:
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}\rightarrow\left|\mathcal{F}\{P\}\right|,$
@@
where $\mathcal{F}\{P\}$ denotes the spatial Fourier transform of the Point Spread Function not affected by the diffuser (as demonstrated in Figure 1d). Note that this is different from direct imaging, because the NA is defined by the combination of the scatterer and the objective [21].
@@
As a result of this analysis, the only information lost is the complex phase of the Fourier transform of the object image without scattering. As demonstrated previously [7, 6, 10], this phase can be retrieved using a phase retrieval algorithm (like in Figure 1e). In our work, we utilized the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms [20].
In practice, we didn’t directly average all the moduli of the Fourier transforms. Instead, we utilized several filtering techniques, such as removing the Gaussian envelope of the image and applying Hann windowing like in [7, 10]. For more details, please refer to the Methods section 5 and Supplemental Material.
@@ 3 Results
To demonstrate the effectiveness of our method in the single photon counting regime, we first conducted a computer simulation. Figure 2 presents the simulation results of imaging the object through the dynamic diffuser in a low photon number regime. Figure 2a shows an object consisting of several point emitters. Imaging this object through an imaging system with finite resolution gives the image shown in Figure 2b ($I=O*PSF$), with the Fourier transform modulus shown in Figure 2c (|$\left|\mathcal{F}\{I\}\right|=\left|\mathcal{F}\{O\}\right|\left|\mathcal{F}\{PSF\}\right|$).
@@
In the simulation, we generated 10,000 frames, each with a different diffuser realization and an average of 400 photons per frame (0.04 photons per pixel). Figure 2f shows the root mean square of the Fourier transform moduli of these 10,000 frames ($\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}$). Notably, Figure 2f closely resembles Figure 2c. After applying the phase retrieval algorithm [20], we obtain the reconstructed image (Figure 2g), which closely resembles the direct image of the object in Figure 2b.
Figure 3: Imaging of Alexa Fluor 568 stained astrocytes through the dynamic complex medium. (a) Schematic of the wide-field fluorescence microscope utilized in our experiments. (b) Direct image of the sample without the diffuser (ground truth). (c) Single frame captured by the qCMOS camera. (d) Average of all 1872 frames recorded. (e) Reconstruction result using the phase retrieval algorithm [20].
In our experiment, we applied our method to fluorescence microscopy. As an experimental setup, we utilized a wide-field custom-made microscope (see Figure 3a). The details of the microscope construction can be found in Section 5. Our microscope allowed for the insertion and removal of a dynamic diffuser, which was placed just before the microscope objective. As a sensor in our microscope, we used the Hamamatsu ORCA-Quest 2 qCMOS camera, which enables single photon counting. We demonstrated our experimental setup on a sample of astrocytes stained with the popular fluorescent dye Alexa Fluor 568. Details of the sample preparation procedure are provided in Methods section 5. Figure 3b shows a direct image of the sample without the diffuser. Figure 3c illustrates a single frame captured by the qCMOS camera. During a frame duration of 200.96 ms, we recorded an average of 1.85 photons per camera pixel. Figure 3d shows the root mean square of all 1872 frames recorded in our experiment. Figure 3d corresponds to the incoherent sums of all frames. Speckles are blurred out and don’t exhibit any spatial information that could be retrieved. Therefore, we average the Fourier transform moduli of all the frames and apply a phase retrieval algorithm [20]. The reconstruction result using the phase retrieval algorithm [20] is presented in Figure 3e.
@@
Figure 4: Imaging of fluorescence micro-spheres through the dynamic complex medium. (a) Direct images of fluorescent microspheres without the diffuser. (b) The modulus of the Fourier transforms of the images in column (a). (c) The average of the Fourier transform moduli of the frames captured through the dynamic diffuser. (d) Recovered image computed using the phase retrieval algorithm [20].
@@
To demonstrate the versatility of our method, we applied it to various objects constructed with fluorescent micro-spheres samples (see Figure 4). The sample preparation procedure for the micro-spheres is detailed in Section 5. The column a of Figure 4 shows direct images of the fluorescent micro-spheres in our microscope. Figure 4b presents the modulus of the Fourier transforms of the images in the column a. Figure 4c shows the average of the Fourier transform modulus of the frames captured by the camera through the dynamic diffuser. Figure 4d displays the recovered image imaged through the diffuser, computed using the column c from fig. 4 with the phase retrieval algorithm [20].

### v2

@@ Abstract
Optical imaging plays a crucial role in advancing science and technology, enabling applications in fields ranging from biomedicine to astronomy. However, imaging through scattering media such as biological tissues, fog, or turbulent atmosphere remains a major challenge. Light scattering and absorption in such media make imaging challenging; in the case of time varying scatterers and low-light regime imaging has not been demonstrated so far. We present the first demonstration of non-invasive imaging of dim objects hidden behind dynamic scattering layers, obtaining robust reconstruction even at extremely low photon counts per frame. We achieve this by developing a new data-processing approach. In our experiment, we utilize a photon number resolving camera to capture a sequence of frames, containing on average, less than one photon per pixel. We validate our approach in microscopy, where we reconstruct images of biological samples stained with standard fluorescent dyes. Beyond microscopy, our approach can be applied in LIDAR systems for imaging through fog, and endoscopy using multimode and multicore fibers.
@@ 1 Introduction
Light scattering plays a crucial role in determining the appearance of most of the things we see. When scattering occurs multiple times within a medium, such as biological tissues, fog, or clouds, it is called a complex medium. The propagation of light through complex media results in random interference patterns known as speckle patterns[1]. For a long time, light scattering in complex media was considered uncontrollable. This fundamental challenge of light scattering affects not only imaging [2], but also poses obstacles in communication, spectroscopy, and other optical and photonic technologies.
Imaging through complex media has long been seen as a formidable challenge. However, a groundbreaking experiment by Vellekoop and Mosk [3] demonstrated that by controlling wavefronts through spatial light modulators (SLMs), light can be focused through scattering media. Subsequent experiments by Popoff et al. significantly advanced the field by introducing a method to measure the transmission matrix of complex media [4, 5]. This development enabled light control behind those media using SLMs.
These remarkable advancements led to a better understanding of complex media and paved the way for developing non-invasive imaging techniques through such media. Specifically, experiments by Bertolotti et al. or Katz et al. [6, 7] utilized the phenomenon known as the memory effect. Within small regions, when an input wavefront reaching a complex medium is tilted within a certain angular range, the output wavefront is equally tilted, resulting in the translation of the speckle pattern behind the complex medium [8, 9]. In imaging, within a memory effect region, light emitted from each point source behind the complex medium creates an almost identical speckle pattern that is shifted according to the position of the source origin.
@@
While several techniques for optical imaging in complex media were developed and represent significant achievements, their widespread adoption remains limited. One limitation is that these techniques typically work with static diffusers. In live bioimaging, biological tissues change over time, necessitating methods that can handle dynamic scattering. Similarly, LIDAR must operate in fog, where light scattering changes over time. Some approaches to overcome this issue involve using high-speed digital micromirror devices (DMDs) [15, 16, 17, 18, 19] as SLMs for fast transmission matrix measurement. However, all these methods require high signal levels, which hinder their practical applications. For rapidly changing diffusers, one must acquire light over a short period, to approximate the diffuser as static. This inevitably results in low light intensities being captured. This creates a growing need for robust imaging techniques that work in low-light intensity scenarios through dynamic complex media.
We demonstrate that non-invasive imaging behind dynamic scattering layers is possible even in the low-photon number regime. There was no previous demonstration under such conditions. To achieve this, we develop a data-analysis approach that reconstructs the object from a sequence of ultra-dim frames recorded behind a scattering layer, without ballistic light. Our imaging method builds on concepts introduced in earlier works [6, 7, 10], but extends their capabilities to enable imaging through dynamically changing diffusers and under low-light conditions.
The proposed approach operates even under extreme cases, only two photons per frame, and remains effective in highly dynamic scenarios, as it does not require the scattering medium to remain static between frames. We present a proof-of-concept experimental demonstration of our method in a widefield microscope imaging fluorescent samples. We use a single-photon camera operating in an ultra-low-light regime, recording fewer than one photon per pixel per frame. Our method does not require wavefront shaping, transmission matrix characterization, or light focusing through the complex medium.
@@ 2 Principles
Figure 1: Concept of the method. (a) A luminescent, spatially incoherent object is imaged through a time-varying diffuser and recorded with a high-speed, photon-counting camera. (b) A long exposure produces a homogeneous image that carries no spatial information about the object. (c) A single short-exposure frame preserves speckle structure but contains very few photons. (d) In the analysis, we use a series of short frames with a low signal. (e) Summing many such low-signal frames similarly yields a homogeneous image without spatial content. (f) Instead, for each short-exposure frame, we compute the magnitude of its Fourier transform. (g) The root-mean-square (RMS) average of these Fourier magnitudes across the burst recovers the spatial-frequency content of the object. (h) After retrieving the missing Fourier phase using a phase-retrieval algorithm [20], the object hidden behind the dynamic diffuser can be reconstructed.
Our approach involves recording a movie of a luminescent object placed behind a changing diffuser (see Fig. 1a). When the camera integration time is long, the diffuser undergoes substantial changes during a single exposure (see Fig. 1b). As a result, the recorded speckle patterns average out, yielding a spatially homogeneous image that carries no information about the hidden object. If the exposure time is short, the frame can be approximated as corresponding to a single realization of the diffuser. However, such frames may contain very low photon counts, sometimes down to a few detected photons per frame (see Fig. 1c).
In our method, we acquire a sequence of short exposure frames (see Fig. 1d). A naive integration of these frames would produce an image similar to that obtained with a single long-exposure measurement (see Fig. 1e), again lacking spatial information. Instead, for each frame we compute the modulus of its two-dimensional Fourier transform (see Fig. 1f). We then average these Fourier-modulus images (see Fig. 1g). Remarkably, the resulting averaged matrix corresponds to the Fourier-transform modulus of the object hidden behind the diffuser and can be obtained even when individual frames contain only single detected photons. With this modulus, the object image can be reconstructed (see Fig. 1h) using well established phase-retrieval algorithms.
This method enables effective imaging under dynamic scattering media conditions and achieves reliable results in low-photon regimes that allow tackling extreme low-light and dynamic scattering conditions, which are completely beyond the reach of currently demonstrated techniques.
Figure 2: Mathematical background of our method. (a) Direct imaging in a scattering-free system. The image $I$ formed on the camera is a convolution of the object $O$ with the system Point Spread Function (PSF) $h$. (b) Under dynamic scattering, each speckle image $I_{i}$ formed on the camera is a convolution of the object $O$ with the instantaneous speckle Point Spread Function (PSF) $S_{i}$. (c) Spatial Fourier transforms of the images shown in (b). (d) The root mean square (RMS) of the Fourier transform moduli, computed across many frames, converges to $|\mathcal{F}\{O\}|\,|\mathcal{F}\{P\}|$, where $P$ denotes the effective PSF of the system as if no scattering were present. (e) The magnitude of the scattering-free Fourier transform is obtained, while its phase is lost. A phase retrieval algorithm [20] is then used to recover the missing phase information. (f) Reconstruction of the object after phase retrieval.
In a conventional, scattering-free imaging system, the image recorded by the camera $I$ is a 2D mathematical convolution of the object $O$ and the system’s Point Spread Function (PSF) $h$ (see Fig. 2a). In the presence of scattering, however, the situation becomes more complex. The pattern formed on the camera, known as a speckle pattern, does not resemble an ordinary PSF. Nevertheless, for small objects located within the memory-effect range [8, 9], each point of the object produces the same speckle pattern, shifted relative to one another. As a result, the image captured by the camera is the convolution of the object with this speckle pattern, which acts as a new speckle Point Spread Function.
Under dynamic scattering conditions, each speckle image formed on camera is a convolution of the object $O$ with the instantaneous speckle Point Spread Function (PSF) $S_{i}$ of the varying diffuser (as illustrated in Figure 2b). We can represent each speckle image as:
$I_{i}=O*S_{i},$ (1)
where $i=1,2,3,\ldots,M$ is the image index, and $M$ is the total number of images. Note that, each frame $F_{i}$ captured by the camera is a Poisson-distributed measurement of the underlying noiseless speckle image.
Under this model, the Fourier transforms satisfy:
$\mathcal{F}\{I_{i}\}=\mathcal{F}\{O\}\mathcal{F}\{S_{i}\},$ (2)
@@
where $\mathcal{F}\{I_{i}\}$, $\mathcal{F}\{O\}$, and $\mathcal{F}\{S_{i}\}$ denote the spatial Fourier transforms of each image $I_{i}$, the object $O$, and the instantaneous speckle Point Spread Function $S_{i}$, respectively (as shown in Fig. 2c). By computing the root mean square of the Fourier transform moduli across all images (Fig. 2d), we obtain:
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}.$ (3)
For many realizations of the diffuser, the root mean square of the Fourier transform moduli of speckle Point Spread Functions $S_{i}$ averages [1] to Fourier transform of the effective Point Spread Function ($\mathcal{F}\{P\}$) of the imaging system as it would not have been affected by the diffuser (see Fig. 2d). Among common averaging strategies, the RMS operator provides the fastest and most reliable convergence to the true scattering-free Fourier magnitude, as demonstrated by our statistical analysis in the Supplementary Information, Section 2.1.2.
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}\;\xrightarrow[M\to\infty]{}\;\left|\mathcal{F}\{P\}\right|.$ (4)
Because each recorded frame $F_{i}$ is a Poisson-distributed measurement of the underlying noiseless speckle image, its Fourier magnitude also satisfies a law-of-large-numbers convergence. Therefore,
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{F_{i}\}\right|^{2}}\;\xrightarrow[M\to\infty]{}\;\left|\mathcal{F}\{O\}\right|\,\left|\mathcal{F}\{P\}\right|,$ (5)
In the Supplementary Information Section 2.1.2, we provide a comprehensive statistical evaluation, estimating the required number of frames for image reconstruction, the achievable resolution, and the signal-to-noise ratio (SNR) of the recovered Fourier magnitude.
@@
Note that in this setting, imaging through the diffuser is different from direct imaging, because the effective numerical aperture NA is defined by the combination of the scatterer and the objective [21].
@@
As a result of this analysis, we retrieve the magnitude of the Fourier transform of the image of the object that is not affected by scattering (see Fig. 2e), and information about its complex phase is lost. As demonstrated previously [7, 6, 10], this phase can be retrieved using a phase retrieval algorithm. In our work, we utilized the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms [20]. Once the Fourier phases are estimated, the object obscured by the dynamic diffuser can be retrieved (like in Figure 2f).
We detail the post-processing techniques and algorithms used in our analysis in Supplementary Information Sections 2.1.4 and 3.
It is important to note that, even though phase retrieval algorithms are highly sensitive to noise, we do not apply them to noisy raw frames. Instead, we apply phase retrieval to the estimated high-SNR Fourier modulus of the object hidden behind a dynamically changing scatterer. For more details of data analysis, please refer to the Methods Section 5 and Supplementary Information Section 3.
@@ 3 Results
To demonstrate the effectiveness of our method in the single photon counting regime, we first conducted a computer simulation. Figure 3 presents the simulation results of imaging the object through the dynamic diffuser in a low photon number regime. Figure 3a shows an object consisting of several point emitters. Imaging this object through an imaging system with finite resolution gives the image shown in Figure 3b (I=OPSFI=O*PSF), with the Fourier transform modulus shown in Figure 3c ($\left|\mathcal{F}\{I\}\right|=\left|\mathcal{F}\{O\}\right|\left|\mathcal{F}\{PSF\}\right|$).
@@
In the simulation, we generated 10,000 frames, each with a different diffuser realization and an average of 400 photons per frame (0.04 photons per pixel). Figure 3f shows the root mean square of the Fourier transform moduli of these 10,000 frames ($\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}$). Notably, Figure 3f shows excellent agreement with Figure 3c. After applying the phase retrieval algorithm [20], we obtain the reconstructed image (Figure 3g), which corresponds to the direct image of the object in Figure 3b.
[…]
In our experiment, we applied our method to fluorescence microscopy. As an experimental setup, we utilized a wide-field custom-made microscope (see Figure 4a). The details of the microscope construction can be found in Section 5. Our microscope allowed for the insertion and removal of a dynamic diffuser, placed just before the microscope objective (between the objective and the fluorescent sample). As a sensor in our microscope, we used the Hamamatsu ORCA-Quest 2 qCMOS camera, offering photon number resolution at each pixel.
[…]
@@
Figure 5: Results comparison of imaging through the dynamic complex medium. (a) Direct images of fluorescent samples without the diffuser. (b) The modulus of the Fourier transforms of the images in column (a). (c) Example camera frames of objects in column (a) obscured by the varying diffuser, with photon counts as low as 0.14 photons per pixel in 1152 frames, highlighted in the second row. (d) The average of the Fourier transform moduli of the frames captured through the dynamic diffuser. (e) Recovered image computed using the phase retrieval algorithm [20]. Rows 1 to 3 present fluorescent microspheres, while row 4 shows Alexa Fluor 568-labeled astrocytes.
@@
To demonstrate the versatility of our method, we applied it to various objects composed of fluorescent micro-spheres and astrocytes stained with the popular fluorescent dye Alexa Fluor 568 (see Figure 5). Fluorescent samples preparation procedures are detailed in Section 5.4. The column a of Figure 5 shows direct images of the fluorescent samples in our microscope. Figure 5b presents the modulus of the Fourier transforms of the images in the column a. Figure 5c presents individual frames for different objects captured through the varying diffuser, with photon counts reaching as low as 0.14 photons per pixel in 1152 captured frames, as shown in the second row. Figure 5d shows the average of the Fourier transform modulus of the frames captured by the camera through the dynamic diffuser. Figure 5e displays the recovered image imaged through the diffuser, computed using the column d from fig. 5 with the phase retrieval algorithm [20]. Rows 1 to 3 present results obtained with fluorescent microspheres, while row 4 shows data from Alexa Fluor 568-labeled astrocytes.